OpenStudy (anonymous):

PLEASE HELP !! A quadratic equation is shown below: 9x^2 - 16x + 60 = 0 Part A: Describe the solution(s) to the equation by just determining the radicand. Show your work. Part B: Solve 4x2 + 8x - 5 = 0 by using an appropriate method. Show the steps of your work, and explain why you chose the method used. Part C: Solve 2x2 -12x + 5 = 0 by using a method different from the one you used in Part B. Show the steps of your work.

OpenStudy (anonymous):

9x^2 - 16x + 60 = 0 comparing it with ax^2 + bx + c = 0 a = 9 b=-16 c=60 b^2 => 256 < 2160 = 4*9*60 Hence the Equation has Complex Roots. 4x2 + 8x - 5 = 0 a = 4 b = 8 c = -5 b^2 = 64 > -80 <= 4*4*-5 Hence Real Roots Exist. sum of Roots (p+q): -2 Product (pq): -5/4 (p-q)2 = (p+q)2 - 4pq p-q = root (4+5) => 3 or -3 p+q = -2 => p=1/2 (+3) or -5/2 (-3) => q=-5/2 (+3) or 1/2 (-3) Roots are 1/2 , -5/2 2x2 -12x + 5 = 0 a = 2 b = -12 c = 5 b2 = 144 > 40 <= 4*2*5 Hence Real Roots Exist The Roots are: -b +/- root(b2-4ac)/2a => 12 +/- root(104)/4 => 12 +/- 2root(26)/4 => 6+root(26)/2 or 6-root(26)/2 Hope its Correct...